Solve the following homogeneous differential equation 2x3 y' = y(2x2− y2


1
Expert's answer
2021-08-26T15:06:22-0400

"\\displaystyle\n2x^3y'\u00a0= y(2x^2 - y^2)\u00a0\\\\\n\n\\textsf{Substitute}\\,\\,\\, y = vx\\\\\n\n2x^3(xv' + v) = 2x^2\\cdot vx - v^3 x^3\\\\\n\n2(xv' + v) = 2v - v^3\\\\\n\n2xv' = -v^3\\\\\n\n2x \\frac{\\mathrm{d}v}{\\mathrm{d}x} = -v^3\\\\\n\n\\frac{\\mathrm{d}v}{v^3} = -\\frac{\\mathrm{d}x}{2x}\\\\\n\n-\\int \\frac{2\\mathrm{d}v}{v^3} = \\int\\frac{\\mathrm{d}x}{x}\\\\\n\n\\frac{1}{v^2} = \\ln{x} + C\\\\\n\nv^2 = \\frac{1}{\\ln{x} + C}\\\\\n\n\\frac{y^2}{x^2} = \\frac{1}{\\ln{x} + C}\\\\\n\n\\implies y^2 = \\frac{x^2}{\\ln{x} + C}"